OpenStudy (anonymous):

What is the domain of the function. y=square root of 3x+3 B. y=square root of 7x-12

OpenStudy (anonymous):

\[y=\sqrt{3x+3}\]

OpenStudy (cwrw238):

Hint for real values of the function 3x+3 cannot be negative so any value of which makes it negative is not in the domain

OpenStudy (cwrw238):

* any value of x which makes it negative

OpenStudy (cwrw238):

note that there is no real square root of a negative

OpenStudy (cwrw238):

so for example x cannot be -2 because that would make 3x + 3 = -3

OpenStudy (anonymous):

b. \[y=\sqrt{7x-12}\]

OpenStudy (anonymous):

dont get it

OpenStudy (anonymous):

@cwrw238

OpenStudy (ranga):

For the first problem, 3x+3 cannot be negative is same as saying \[3x + 3 \ge 0\]. Solve for x.

OpenStudy (anonymous):

oh so it would be

OpenStudy (anonymous):

\[x \le-1\]

OpenStudy (anonymous):

@ranga

OpenStudy (anonymous):

and B would be \[x \ge-\frac{ 7 }{ 12}\]